Federal contractors with contracts in excess of $50,000 and 50 or more employees are subject to Executive Order 11246 which requires Affirmative Action Plans. As part of their compliance with AA, contractors must make a good faith effort to achieve full utilization of women and minorities in job groups where underutilization exists. Utilization analysis is a comparison of (1) the percentage of minority or women employees employed by an organization in a job group and (2) the percentage of minorities and women available in the relevant labor market in those job groups. The relevant labor market is where the contractor typically recruits or selects members of the job group or labor markets from which such members could reasonably be recruited. If underutilization exists, a goal must be set to move toward full utilization.

1. Contractor A selects engineering supervisors solely from within the organization. Supervisors are selected only from the Project Engineer rank. None are recruited externally. The contractor has 10 engineering supervisors of whom 2 are minorities; none are women. The contractor has 100 project engineers, 10 are minorities and 10 are women. Does underutilization exist? If so, where? If so, what action should the contractor take?

2. Hospital B has Federal research contracts and employs 100 nursing supervisors of which 40 are minorities and 90 are women. The hospital selects approximately 80% of their nursing supervisors from among their own nurses and 20% from the external labor market. The hospital’s nursing job group consists of 60% minorities and 90% women. The regional external labor market for nurses consists of 40% minorities and 90% women. Does underutilization exist? If so, where? If so, what action should the hospital take?

Please show all work (how answers were arrived at).
Answer #1

1. As per definition of underutilization, the percentage of minority or women employees employed by the organization as Engineering Supervisors = (2/10)*100 = 20%

Also, the percentage of minorities available in the relevant labor market in internal Project Engineer rank = 10/100*100 = 10%.

Also, the percentage of women available in the relevant labor market in internal Project Engineer rank = 10/100*100 = 10%.

Since we have more minorities and no women are employed. So underutilization exists, and the organization should employ 10% of 10 viz. 1 woman engineering supervisor as well to move towards full utilization.

2. Yes, underutilization exists at Hospital B and hospital should recruit an additional 16 minority employees from the internal resource pool. Below is the screenshot of calculation in excel along with respective formulas used.

20 Employed By Hospital %ages Internal Employees External Employees Total Employees Minorities 40 40% 40 Women 90 90% 18 90 T

Formulas used -

10 90 100 20 Employed By Hospital %ages Internal Employees External Employees Total Employees Minorities =C3/$C$5 =C3*$E$6 =C